OpenStudy (kamille):

Avogadro’s number NA is 6.023 * 10^23. A sandpit contains 3.0115 billion sand grains. This corresponds to what fraction of NA? How do I think: Since I need to find a fraction, I can make a proportion: Avagardo number/1=number of saind grains/x Do I think correctly?

OpenStudy (kamille):

@phi

OpenStudy (phi):

if the question, was "1 is what fraction of 10?" do you see the answer is 1/10 3 is what fraction of 6: 3/6 = ½

OpenStudy (kamille):

so I need to ignore the 10^12 and 10^23? if yes, how to write this correctly 'in math language'?

OpenStudy (phi):

If we use your idea \[ \frac{NA}{1}= \frac{sand}{x} \] and solve for x, we get \[ x = \frac{sand}{NA} \] which is the same idea... you want \[ \frac{3.0115 \ billion}{6.023 \cdot 10^{23} }\]

OpenStudy (kamille):

I still don't get where 10^12 and 10^23 disappear : (

OpenStudy (phi):

they don't disappear. billion is another name for \( 10^9\) see https://en.wikipedia.org/wiki/Names_of_large_numbers that means your fraction is \[ \frac{3.0115 \cdot 10^9}{6.023 \cdot 10^{23} } \]

OpenStudy (kamille):

yeah, but I get 1/2 *10-14 is this an answer?

OpenStudy (phi):

10^9 is 10*10*10*10*10*10*10*10*10*10= 1000000000 when you divide it by 10^23 (10 with 23 zeros) you get 10^14 in the bottom \[ \frac{3.0115}{6.023 \cdot 10^{14}}\]

OpenStudy (kamille):

or I think correctly it should be 2*10^13, but is this a fraction?

OpenStudy (kamille):

and thanks;))

OpenStudy (phi):

yes you get 0.5 * 10^-14 that is not in standard form (you want the lead number to be between 1 and 10) so multiply the 0.5 * 10 and divide the 10^-14 by 10 (that keeps things balanced) you get 5 * 10^-15

OpenStudy (kamille):

thanks: ))

OpenStudy (kamille):

and are you sure it is 5, not 2?

OpenStudy (phi):

we start with \[ \frac{3.0115 \ billion}{6.023 \cdot 10^{23} } \] \[ \frac{3.0115 \cdot 10^9}{6.023 \cdot 10^{23} } \] \[ \frac{3.0115}{6.023 \cdot 10^{14} } \] \[ \frac{3.0115}{6.023} \cdot 10^{-14} \] the 3.0115/6.023 = ½ \[ \frac{1}{2} \cdot 10^{-14} \] change ½ to a decimal 0.5 \[ 0.5 \cdot 10^{-14} \] now adjust 0.5 by multiplying by 10 (and divide by 10) \[ 0.5 \cdot 10 \cdot 10^{-1} \cdot 10^{-14} \] simplify that to \[ 5.0 \cdto 10^{-15} \]

OpenStudy (phi):

*\[ 5.0\cdot 10{−15} \]

OpenStudy (kamille):

okay, why do we need to change from 1/2 to decimal form?

OpenStudy (phi):

you can write numbers lots of different ways, but math is confusing enough if people don't agree on a few standards.

OpenStudy (kamille):

and we can't simply just write 1/2 as 2, becase you need to change spmething with 10^-14, yeah?

OpenStudy (phi):

you agree ½ of a pie is not the same as 2 pies ? in other words ½ and 2 are not the same thing. however, ½ and 0.5 are the same number. (as a fraction and as a decimal)

OpenStudy (kamille):

ywah, finally I DO understand: )) thanks again

OpenStudy (phi):

but once you write the number as 0.5 * 10^-14 you are still not quite done... 0.5*10 = 5.0 and 10^-14 / 10 = 10^-15 and the answer in standard form is \[ 5.0 \cdot 10^{-15} \]